x^2/16+y^2/9=1
What is the foci and vertices of this ellipse equatioin

Answer by lwsshak3(11628)

x^2/16+y^2/9=1
What is the foci and vertices of this ellipse equation
**
Given equation is that of an ellipse with horizontal major axis of the standard form:
(x-h)^2/a^2+(y-k)^2/b^2=1,a>b, (h,k)=(x,y) coordinates of center.
center:0,0)
a^2=16
a=√16=4
vertices:(0±a,0)=(0±4,0)=(-4,0) and ((4,0)
b^2=9
b=√9=3
c^2=a^2-b^2=16-9=7
c=√7
Foci: (0±c,0)=(0±√7,0)=(-√7,0) and (√7,0)